# The empirical median for estimating the common mean of heteroscedastic random variables

### v1

@@ 4.1 Proof of Theorem 1
The objective is now to bound from below $B(t)$. We set $k_{t}=\sup{(i:t>\sigma_{i})}$. We fix $C$ a positive constant such that $10<C<$ 1. By assumption of the theorem, for all $i\in[\![1,n]\!]$ there exists $\sigma_{i}$ such that if $0\leqslant t\leqslant\sigma_{i}$ then $\mathbb{P}(0\leqslant X_{i}\leqslant t)\geqslant C\frac{t}{\sigma_{i}}$. Then,
@@ 4.2 Proof of Theorem 2
The corollary follows immediately from Lemma 1 and the fact that under assumption of Corollary 2 we have that for all $i\in[\![1,n]\!]$, $p_{i}(t)\in[\frac{1}{4},\frac{3}{4}]$.
We want to bound from below the probability
@@
If $i\leqslant k_{t}$, which is equivalent to $t>\sigma_{i}$, it follows that $\mathbb{P}(X_{i}>t)=\mathbb{P}(X>\frac{t}{\sigma_{i}})\leqslant\frac{1}{4}$. Thus, by Hoeffding’s inequality applied to the first sum, we have
@@
with $c-\frac{1}{\sqrt{2}}\leqslant 0.3$ and $\delta\leqslant\frac{1}{4}$ such that $\log{(\frac{1}{2\delta})}\geqslant\frac{1}{2}\log{(\frac{1}{\delta})}$. So with probability at least $\delta$ we have that
@@
$k_{t}\leqslant\frac{n}{2}\quad\text{and }\quad k_{t}\leqslant\frac{n(\sqrt{2}-1)^{2}}{8}$ (6)
@@
with $k_{t}=\left\lfloor\frac{\sqrt{2}-1}{8}\sqrt{n\log{(\frac{1}{\delta})}}\right\rfloor$.
@@ 4.3 Proof of Lemma 1
Let $t>0$ and $S=\sum_{i=1}^{n}V_{i}$ with $V_{i}\sim\mathcal{B}({p_{i}})$, $i\in[\![1,n]\!]$. We consider a change of measure meaning that we move from $S$, a sum of random variables with Bernoulli distribution of parameter $p_{i}$ to $\tilde{S}$, a sum of random variables with Bernoulli distribution of parameter $\tilde{p_{i}}$ such that
@@
We assume that for all $i\in[\![1,n]\!]$, $p_{i}\in[\frac{1}{4},\frac{3}{4}]$. Let us show that $\tilde{p_{i}}$ is also in the interval $\left[\frac{1}{4},\frac{3}{4}\right]$ for all $i\in[\![1,n]\!]$. We know that $\tilde{p_{i}}=p_{i}+\Delta$, so $\tilde{p_{i}}\geqslant\frac{1}{4}.$ And since $p_{i}\leqslant\frac{1}{2}$, we deduce that $\tilde{p_{i}}\leqslant\frac{3}{4}$ as long as $\Delta<\frac{1}{4}$ which is equivalent to $c<\frac{1}{4}\sqrt{\frac{n}{\log{(\frac{1}{\delta})}}}$ that is verified for $n$ sufficiently large. Since $p_{i}$ and $\tilde{p_{i}}$ are of order $\frac{1}{2}$, the Hoeffding bound is optimal. Now, let $c^{\prime}\in\mathbb{R}$ and we want to bound from below
@@
For that, we need to bound from below $\mathbb{P}\left(\tilde{S}\geqslant\mathbb{E}[S]+c^{\prime}\sqrt{n\log{\left(\frac{1}{\delta}\right)}}\right),$ and to bound from above $\lvert\mathbb{P}(\tilde{S}\geqslant t)-\mathbb{P}(S\geqslant t)\rvert.$
We start by reducing from below $\mathbb{P}\left(\tilde{S}\geqslant\mathbb{E}[S]+c^{\prime}\sqrt{n\log{\left(\frac{1}{\delta}\right)}}\right).$ We have that
@@
We have that
$\mathbb{P}\left(\tilde{S}-\mathbb{E}[\tilde{S}]\leqslant-c^{\prime\prime}\sqrt{n\log{\left(\frac{1}{\delta}\right)}}\right)=\mathbb{P}\left(-\tilde{S}-\mathbb{E}[-\tilde{S}]\geqslant c^{\prime\prime}\sqrt{n\log{\left(\frac{1}{\delta}\right)}}\right).$
Moreover, $-\tilde{S}=\sum_{i=1}^{n}(-\tilde{V_{i}})$ where $(-\tilde{V_{i}})$ are $\frac{1}{8}$-Sub-Gaussian random variables. By Hoeffding’s inequality, we then have that for all $t>0$,
@@
$\tilde{S}=\Phi(Y_{1},\dots,Y_{n})=Y_{1}+\dots+Y_{n},\quad Y_{i}\sim\mathcal{B}(\tilde{{p_{i}}}).$
@@
By [19, (2.25) p.89], we have that $D_{\mathrm{TV}}(P,Q)\leqslant 1-\frac{1}{2}\exp{(-\mathrm{KL}({P},{Q}))}$ and by Lemma $2.7$ in [19], we have that $\mathrm{KL}({P},{Q})\leqslant\log{(1+\chi^{2}(P,Q))}$ where $\chi^{2}$ denotes the $\chi^{2}$ divergence. Besides, for $P_{i},Q_{i},i\in 1,\dots,n$ laws on $E_{i}$ and for $P=\bigotimes_{i=1}^{n}P_{i}$ and $Q=\bigotimes_{i=1}^{n}Q_{i}$ product laws on $\prod_{i=1}^{n}E_{i}$, we have that $1+\chi^{2}(P,Q)=\prod_{i=1}^{n}(1+\chi^{2}(P_{i},Q_{i}))$. Thus,
@@ 4.4 Elementary lemmas and proofs
In this section, we provide for the sake of completeness statements and short proofs of facts used.
@@
we have that $p_{i}(t)\geqslant\frac{1}{4}$ for all $i\in[\![1,n]\!]$.
@@ Acknowledgments
I would like to express my gratitude to my two PhD supervisors for their contribution to the realization of this work. I am indebted to Jaouad Mourtada for his crucial assistance in conceiving the ideas and elaborating the proofs, and to Alexandre Tsybakov for meticulously proofreading the manuscript and offering invaluable advice.

### v3

@@ 4.1 Proof of Theorem 1
The objective is now to bound from below $B(t)$. We set $k_{t}=\sup{(i:t>\sigma_{i})}$. We fix $C$ a positive constant such that $0<C<1$. By assumption of the theorem, for all $i\in[\![1,n]\!]$ there exists $\sigma_{i}$ such that if $0\leqslant t\leqslant\sigma_{i}$ then $\mathbb{P}(0\leqslant X_{i}\leqslant t)\geqslant C\frac{t}{\sigma_{i}}$. Then,
@@ 4.2 Proof of Theorem 2
The corollary follows immediately from Lemma 1 and the fact that under assumption of Corollary 2 we have that for all $i\in[\![1,n]\!]$, $p_{i}(t)\in[\frac{1}{4},\frac{3}{4}]$. We want to bound from below the probability
@@
If $i\leqslant k_{t}$, which is equivalent to $t>\sigma_{i}$, it follows that $\mathbb{P}(X_{i}>t)=\mathbb{P}(\frac{X_{i}}{\sigma_{i}}>\frac{t}{\sigma_{i}})\leqslant\frac{1}{4}$. Thus, by Hoeffding’s inequality applied to the first sum, we have
@@
with $c-\frac{1}{\sqrt{2}}\leqslant 0.3$ and $\delta\leqslant\frac{1}{4}$ such that $\log{(\frac{1}{2\delta})}\geqslant\frac{1}{2}\log{(\frac{1}{\delta})}$. So, with probability at least $\delta$, we have that
@@
$k_{t}\leqslant\frac{n}{2}\quad\text{and }\quad k_{t}\leqslant\frac{n(\sqrt{2}-1)^{2}}{8},$ (6)
@@
with $k_{t}=\left\lfloor\frac{\sqrt{2}-1}{8}\sqrt{n\log{(\frac{1}{\delta})}}\right\rfloor$. ∎
@@ 4.3 Proof of Lemma 1
Proof.
Let $t>0$ and $S=\sum_{i=1}^{n}V_{i}$ with $V_{i}\sim\mathcal{B}({p_{i}})$, for $i\in[\![1,n]\!]$. We consider a change of measure meaning that we move from $S$, a sum of random variables with Bernoulli distribution of parameter $p_{i}$ to $\tilde{S}$, a sum of random variables with Bernoulli distribution of parameter $\tilde{p_{i}}$ such that
@@
We assume that for all $i\in[\![1,n]\!]$, $p_{i}\in(\frac{1}{4},\frac{3}{4})$. Let us show that $\tilde{p_{i}}$ is also in the interval $\left(\frac{1}{4},\frac{3}{4}\right)$ for all $i\in[\![1,n]\!]$. We know that $\tilde{p_{i}}=p_{i}+\Delta$, so $\tilde{p_{i}}>\frac{1}{4}.$ And since $p_{i}<\frac{3}{4}$, there exists $\epsilon$ such that $p_{i}=\frac{3}{4}-\epsilon$. We deduce that $\tilde{p_{i}}<\frac{3}{4}$ as long as $c<\varepsilon\sqrt{\frac{n}{\log{(\frac{1}{\delta})}}}$ that is verified for $n$ sufficiently large. Since $p_{i}$ and $\tilde{p_{i}}$ are of order $\frac{1}{2}$, the Hoeffding bound is optimal. Now, let $c^{\prime}\in\mathbb{R}$ and we want to bound from below
@@
For that, we need to bound from below $\mathbb{P}\left(\tilde{S}\geqslant\mathbb{E}[S]+c^{\prime}\sqrt{n\log{\left(\frac{1}{\delta}\right)}}\right),$ and to bound from above $\lvert\mathbb{P}(\tilde{S}\geqslant t)-\mathbb{P}(S\geqslant t)\rvert$. We start by reducing from below $\mathbb{P}\left(\tilde{S}\geqslant\mathbb{E}[S]+c^{\prime}\sqrt{n\log{\left(\frac{1}{\delta}\right)}}\right).$ We have that
@@
[…]
@@
[…]
@@
By [24, (2.25) p.89], we have that $D_{\mathrm{TV}}(P,Q)\leqslant 1-\frac{1}{2}\exp{(-\mathrm{KL}({P},{Q}))}$ and by Lemma $2.7$ in [24], we have that $\mathrm{KL}({P},{Q})\leqslant\log{(1+\chi^{2}(P,Q))}$ where $\chi^{2}$ denotes the $\chi^{2}$ divergence. Besides, for $P_{i},Q_{i},i\in 1,\dots,n$ laws on $E_{i}$ and for $P=\bigotimes_{i=1}^{n}P_{i}$ and $Q=\bigotimes_{i=1}^{n}Q_{i}$ product laws on $\prod_{i=1}^{n}E_{i}$, we have that $1+\chi^{2}(P,Q)=\prod_{i=1}^{n}(1+\chi^{2}(P_{i},Q_{i}))$. Thus,
@@ 4.4 Elementary lemmas and proofs
In this section, we provide for the sake of completeness statements and short proofs of the facts used.
@@
we have that $p_{i}(t)\geqslant\frac{1}{4}$ for all $i\in[\![1,n]\!]$. ∎
@@ Acknowledgments
I would like to express my gratitude to my two PhD supervisors for their contribution to the realisation of this work. I am indebted to Jaouad Mourtada for his crucial assistance in conceiving the ideas and elaborating the proofs, and to Alexandre Tsybakov for meticulously proofreading the manuscript and offering invaluable advice.